OpenStudy (anonymous):

7 - 9x - (x + 12) < or equal to 25?

OpenStudy (anonymous):

The question is unclear

OpenStudy (anonymous):

how do you solve for x?

OpenStudy (anonymous):

ohk i get it

OpenStudy (anonymous):

1 min

OpenStudy (anonymous):

7-9x-x-12<=23 -5+-10x<=23 -10x<=23+5 -10x<=28 x<=28/-10 x<=-2.8 When u divide by a negative quantity the signs change